# ICSVR: Investigating Compositional and Semantic Understanding in Video Retrieval Models

### v1

@@ Abstract
Video retrieval (VR) involves retrieving the ground truth video from the video database given a text caption or vice-versa. The two important components of compositionality: objects & attributes and actions are joined using correct semantics to form a proper text query. These components (objects & attributes, actions and semantics) each play an important role to help distinguish among videos and retrieve the correct ground truth video. However, it is unclear what is the effect of these components on the video retrieval performance. We therefore, conduct a systematic study to evaluate the compositional and semantic understanding of video retrieval models on standard benchmarks such as MSRVTT, MSVD and DIDEMO. The study is performed on two categories of video retrieval models: (i) which are pre-trained on video-text pairs and fine-tuned on downstream video retrieval datasets (Eg. Frozen-in-Time, Violet, MCQ etc.) (ii) which adapt pre-trained image-text representations like CLIP for video retrieval (Eg. CLIP4Clip, XCLIP, CLIP2Video etc.). Our experiments reveal that actions and semantics play a minor role compared to objects & attributes in video understanding. Moreover, video retrieval models that use pre-trained image-text representations (CLIP) have better semantic and compositional understanding as compared to models pre-trained on video-text data.
@@ 1 Introduction
When compared to images, measuring compositionality is a lot harder in videos. There are multiple reasons to this: First, videos are made-up of time-series image frames with multiple objects & attributes and actions unlike images. Therefore, methods like creating negative pairs, mismatching pairs etc. used for evaluating compositionality in image-language models have very limited scope. Second, even though tasks based on video question answering (VQA)  [25, 6] have been proposed to measure the compositionality, recent studies  [24, 45, 4, 10, 14] have shown that these datasets exhibit single frame bias.
Most of the previous works  [51, 42, 22] focus on understanding the compositionality of image-text models. It mainly involves experimenting with objects & attributes in the text captions and retrieving the images. However, actions play a crucial role in when retrieving videos using text captions. Another important aspect which is often overlooked in the previous studies is the semantics. For example consider the query a guy wearing a red shirt drives a car while talking, the objects & attributes are guy, red shirt and car, the actions are wearing, driving and talking and rest of the words (a, while) form the semantics of the text captions. The video retrieval models can comprehend such queries because of the accurate semantics and compositionality (objects & attributes and actions).
@@
Now consider the following scenarios of the text captions in which (i) objects & attributes are missing (a wearing a drives a while talking) (ii) actions are missing (a guy a red shirt a car while) and (iii) semantics are missing (guy wearing red shirt drives car talking). This begs an important question: What is the effect of each of these scenarios on the video retrieval performance? To address this question, we propose a detailed study to evaluate the semantic and compositional understanding of video retrieval models.
@@
For this study we create a comprehensive test bed to evaluate the state-of-the-art video retrieval models for compositonality and semantics. We base this investigation along three axes: Objects & attributes, actions and semantics. We propose a set of 10 tasks for these categories: four tasks to evaluate the knowledge of VR models for objects & attributes (§3.1.1), three tasks for testing action understanding (§3.1.2) and finally, three tasks for semantic capabilities (§3.2). Table  1 describes these tasks with an example. We perform a comprehensive evaluation on 12 state-of-the-art video retrieval models belonging to two categories (§4.1): The first category of models such as Frozen-in-Time (FiT)  [3], MCQ  [16] etc. are pre-trained on large scale video datasets and fine-tuned for video retrieval. The second category uses pretrained image features like CLIP for video retrieval namely CLIP4Clip  [35], CLIP2Video  [11] etc. These models are tested on three standard video retrieval benchmarks (§4.2) such as MSRVTT  [55], MSVD  [5] and DiDeMo  [2].
@@
Our experiments (§5.1) reveal that objects & attributes are the most crucial to video retrieval followed by actions and semantics. Among video retrieval models, CLIP based models have a better compositional and semantic understanding when compared with pretrained video models. We further perform detailed studies to fully judge how retrieval models perceive each of the components. We find that (§5.2) video retrieval models have a poor understanding of relationship between objects and its attributes. However, they are extremely sensitive to incorrect object references in the captions. Our studies on action understanding (§5.3) disclose that models have poor sense of action negation and replacing them with incorrect actions lead to slight decrease in video retrieval performance. Finally, we discover (§5.4) that models perform significantly better even without the right semantic word order. Our contributions are as follows:
@@
  • Ours is the first work to comprehensively investigate the compositional and semantic understanding of video retrieval models and we propose a set of 10 tasks for this study.
@@ 2.1 Video retrieval
In a parallel direction, image features pre-trained on large amounts of image-text data have been adopted for the task of video retrieval. CLIP4Clip  [35] is an end-to-end trainable video retrieval model based on CLIP  [44] architecture in which frame features are extracted using clip image encoder and the temporal modelling is performed using a transformer encoder. A two-stage framework CLIP2Video  [11] is proposed to enhance interaction among video features and video-text features for video-retrieval. Madasu et at.  [37] used off-the-shelf multi-lingual data to enhance the performance of video-retrieval. All these video-retrieval models previously haven’t been tested for semantic and compositional understanding. To the best of our knowledge, ours is the first work to comprehensively explore semantic and compositional understanding of video retrieval models.
2.2 Semantics
Transformer based language models  [8, 21, 56] have achieved state-of-the-art results on most natural language understanding tasks  [54, 53]. Hence, there has been a growing interest to explore the semantic capabilities of these models  [58, 46, 52, 43, 15]. More recently, a body of works  [49, 23, 1, 38] investigated the effect of word order and its effect on the semantic understanding in pre-trained language models. Since all the video retrieval models use pre-trained language models for encoding text captions, we build upon those works and investigate their semantic understanding.
@@ 2.3 Compositionality
Although vision-language models pretrained on large amounts of data achieved state-of-the-results there has been a growing interest to understand the working of these models  [29, 25, 6, 41, 25, 50]. These works mainly focus on compositional knowledge these models by proposing new benchmarks. Winoground  [51] dataset was introduced in which a pair of text captions contain the same set of words but pertain to different images. The models are then tested for image and caption match. Another benchmark CREPE  [14] was put forward to evaluate two aspects of compositionality: systematicity and productivity. This benchmark contains unseen compounds and atoms in the test split to evaluate the models’ generalization. Parcalabescu et al.  [42] proposed VALSE dataset to measure visio-linguistic capabilities of pretrained vision and language models. AGQA-Decomp  [14] is a new benchmark to measure compostional consistency for the task of Video Question Answering. All these works proposed new benchmarks for compostional reasoning in image-language models. Contrary to these, our work focuses on measuring compositionality of video retrieval models using the standard datasets and doesn’t require a new benchmark. Moreover, our experiments are evaluated on 13 models which are significantly higher than the models in these works.
@@
$Q_{semrem}$ Semantics removal guy wearing red shirt drives car talking
@@
3 Compositional and Semantic Understanding
In this section, we first define semantics and compositionality and subsequently establish the evaluation protocol for semantic and compositional understanding in video retrieval models. For this evaluation, we augment the existing text captions and create new datasets that assess their semantic and compositional understanding. We explain this protocol using an example test caption $Q$ ”a guy wearing a red shirt drives a car while talking“ from the MSRVTT dataset. Table  1 summarizes different augmentation methods used for the proposed study.
@@ 3.1.1 Object & Attribute knowledge
Object & Attribute removal ($Q_{objattrrem}$):
In this setup, we remove all the objects & attributes in the original caption $Q$ and the resulting caption is ”wearing a drives a while talking“. Here guy, red shirt and car are the objects & attributes.
Object shift ($Q_{objshift}$):
To test the VR models ability to relate objects with their attributes, we shift the places of objects in the captions. The modified caption is ”a shirt wearing a red car drives a guy while talking“.
Object replacement ($Q_{objrep}$):
We evaluate the VR models sensitivity to objects by randomly replacing the objects with an entirely different objects. The replaced caption is ”a surf wearing a red mars drives a guy channel while talking“.
Object partial ($Q_{objpartial}$):
In this setup, the VR models are given access to just 50% of the objects in the caption. This is to understand if the models perform any shortcuts while retrieving videos. Eg: ”a wearing a red drives a car while talking“. Next, we introduce the tasks for evaluating action knowledge in VR models.
@@ 3.1.2 Action knowledge
Another important compositional component in the captions is action. To test the action knowledge of video retrieval models, we propose the following caption augmentations:
Action removal ($Q_{actrrem}$): The actions present in the original captions are eliminated. The modified caption is ”a red shirt a car while“ as the actions wearing, drives are removed. This is to understand the influence of actions on the video retrieval performance.
Action negation ($Q_{actneg}$): A negation is added to all the actions in the captions resulting in the new caption ”a guy not wearing a red shirt not drives a car while not talking“ This tests the VR models ability to comprehend negation in the captions.
Action replacement ($Q_{actrep}$): In this setup, the actions are randomly replaced with a different set of actions. The replaced actions are neither antonyms nor synonyms. It checks if the models truly recognize the meaning of the action words. Next we present the evaluation protocol for semantic understanding of VR models.
3.2 Semantic understanding
In the previous section we elucidated the components for compositional reasoning in videos namely objects & attributes and actions. These components are bind together by semantics there by forming a meaningful caption. Let’s consider a part of the example described previously ”a guy wearing a red shirt drives a car”, if the word ”car“ and ”guy“ are interchanged the resulting caption will be ”a car wearing a red shirt drives a guy” which is not meaningful. Consequently, semantics also play a crucial role in video retrieval performance along with the compositonality. Subsequently, we put forward the evaluation protocol to measure semantic understanding in video retrieval models.
Semantics removal ($Q_{semrem})$: Our first experiment focuses on the effect of semantics on VR models. We modify the caption by keeping just the objects & attributes, actions and eliminate any semantic words among them. The resulting caption is ”guy wearing red shirt drives car talking“.
Word order shuffle ($Q_{shuf})$: In this setup, all the words are shuffled in the caption. This destroys the order of compositionality and semantics. This tests the order sensitivity of VR models.
Word order reverse ($Q_{rev})$: In this setup, we preserve the word order except that in the reverse order. It evaluates the positional knowledge of video retrieval models. Next, we present the experiment set up for quantifying the compositional and semantic understanding.
@@ 3.2 Semantic understanding
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$
@@
Table 2: The table shows the results on MSRVTT  [55] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{semrem}$ is the R@1 score on datasets that have excluded actions, attributes and semantics respectively.
@@
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$
@@
Table 3: The table shows the results on MSVD  [5] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{semrem}$ is the R@1 score on datasets that have excluded actions, attributes and semantics respectively.
@@
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{semrem}$
@@
Table 4: The table shows the results on DiDeMo  [2] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{semrem}$ is the R@1 score on datasets that have excluded actions, attributes and semantics respectively.
@@ 4.3 Implementation
We use spacy [^1] https://spacy.io/usage/linguistic-features to identify parts-of-speech for all the words in the caption. We consider nouns, adverbs and adjectives as objects & attributes, verbs as actions and rest of the parts-of-speech as semantics. We use the exact set up used by the state-of-the-art video retrieval models and measure the performance on all the augmented datasets.
@@ 5 Results and Discussion
5.1 Objects & Attributes vs Actions vs Semantics: Do all of them matter?
Our aim is to analyze the importance of three components: objects & attributes, actions and semantics that make up a text query for retrieving videos. Hence, we test the video retrieval models with text captions that have missing objects & attributes ($Q_{objattrrem}$), actions ($Q_{actrem}$) and semantics ($Q_{semrem}$). Tables  2,  3 and  4 show the results on MSRVTT, MSVD and DiDeMo datasets respectively.
@@ 5.1 Objects & Attributes vs Actions vs Semantics: Do all of them matter?
It is evident from the table that there is a drop in video retrieval performance when tested with text captions that don’t have actions ($Q_{actrem}$). The drop is more pronounced among CLIP based models than pretrained video models. This shows that actions play a role for retrieving correct videos. However, we see that the performance drop is not as expected. Videos are time-series image frames which can have same attributes. In those scenarios, actions help in differentiating those videos. We see this effect when the R@1 is lower among pretrained video models and higher among CLIP based models. When the performance is lower, actions do not play a significant role in video retrieval and hence the videos can be retrieved without them in the text caption. On the contrary if R@1 score is higher, we see a notable decline. This is due to the robust video representations of CLIP based models as compared to pretrained video models. CLIP based models accurately encode video representations but, when the differentiating factor among videos i.e actions are missing in text captions leads to retrieval of incorrect videos. In short caption length datasets like MSRVTT and MSVD, we notice a significant drop in performance as compared with DiDeMo which is a paragraph ($>$ 1 sentences) dataset. This is because text captions in DiDeMo contains detailed description of the videos and hence, missing actions didn’t lead to drop in performance as compared to MSRVTT and MSVD. It demonstrates that actions are not essential in paragraph-video retrieval.
@@
Next, we analyze the performance of video retrieval models tested with text captions without semantics ($Q_{semrem}$). From the table, it is clear that there is a reduction in R@1 without the semantics in the text captions. It validates that semantics are necessary for retrieving correct ground truth videos. For MSRVTT, we observe that models tested without semantics under-perform compared to actions in the text captions and the average difference in performance is 2%. The reverse is true for MSVD and DiDeMo datasets where there is a huge difference of 9%. In addition, we also notice that CLIP based models are more sensitive to semantics than pretrained video models. Finally, we evaluate the video retrieval models on text captions in the absence of objects & attributes ($Q_{objattrrem}$). As seen from the results, these models perform poorly (a drop in 20%) which underscores the significance of objects & attributes. We also notice that $Q_{attrrem}$ trails $Q_{actrem}$ and $Q_{semrem}$ by a huge margin. This difference is more striking among CLIP based models as opposed to pretrained video models.

### v3

# ICSVR: Investigating Compositional and Syntactic Understanding in Video Retrieval Models

@@ Abstract
Video retrieval (VR) involves retrieving the ground truth video from the video database given a text caption or vice-versa. The two important components of compositionality: objects & attributes and actions are joined using correct syntax to form a proper text query. These components (objects & attributes, actions and syntax) each play an important role to help distinguish among videos and retrieve the correct ground truth video. However, it is unclear what is the effect of these components on the video retrieval performance. We therefore, conduct a systematic study to evaluate the compositional and syntactic understanding of video retrieval models on standard benchmarks such as MSRVTT, MSVD and DIDEMO. The study is performed on two categories of video retrieval models: (i) which are pre-trained on video-text pairs and fine-tuned on downstream video retrieval datasets (Eg. Frozen-in-Time, Violet, MCQ etc.) (ii) which adapt pre-trained image-text representations like CLIP for video retrieval (Eg. CLIP4Clip, XCLIP, CLIP2Video etc.). Our experiments reveal that actions and syntax play a minor role compared to objects & attributes in video understanding. Moreover, video retrieval models that use pre-trained image-text representations (CLIP) have better syntactic and compositional understanding as compared to models pre-trained on video-text data. The code is available at https://github.com/IntelLabs/multimodal_cognitive_ai/tree/main/ICSVR.
@@ 1 Introduction
When compared to images, measuring compositionality is a lot harder in videos. There are multiple reasons to this: First, videos are made-up of time-series image frames with multiple objects & attributes and actions unlike images. Therefore, methods like creating negative pairs, mismatching pairs etc. used for evaluating compositionality in image-language models have very limited scope. Second, even though tasks based on video question answering (VQA) [23, 6] have been proposed to measure the compositionality, recent studies [22, 42, 3, 9, 13] have shown that these datasets exhibit single frame bias. Most of the previous works [48, 39, 21] focus on understanding the compositionality of image-text models. It mainly involves experimenting with objects & attributes in the text captions and retrieving the images. However, actions play a crucial role in when retrieving videos using text captions. Another important aspect which is often overlooked in the previous studies is the syntactics. For example consider the query a guy wearing a red shirt drives a car while talking, the objects & attributes are guy, red shirt and car, the actions are wearing, driving and talking and rest of the words (a, while) form the syntactics of the text captions. The video retrieval models can comprehend such queries because of the accurate syntactic and compositionality (objects & attributes and actions).
@@
Now consider the following scenarios of the text captions in which (i) objects & attributes are missing (a wearing a drives a while talking) (ii) actions are missing (a guy a red shirt a car while) and (iii) syntactics are missing (guy wearing red shirt drives car talking). This begs an important question: What is the effect of each of these scenarios on the video retrieval performance?
@@
To address this question, we propose a detailed study to evaluate the syntax and compositional understanding of video retrieval models. For this study we create a comprehensive test bed to evaluate the state-of-the-art video retrieval models for compositonality and syntactics. We base this investigation along three axes: Objects & attributes, actions and syntactics. We propose a set of 10 tasks for these categories: four tasks to evaluate the knowledge of VR models for objects & attributes (§3.1.1), three tasks for testing action understanding (§3.1.2) and finally, three tasks for syntactic capabilities (§3.2). Table  1 describes these tasks with an example. We perform a comprehensive evaluation on 12 state-of-the-art video retrieval models belonging to two categories (§4.1): The first category of models such as Frozen-in-Time (FiT) [2], MCQ [15] etc. are pre-trained on large scale video datasets and fine-tuned for video retrieval. The second category uses pretrained image features like CLIP for video retrieval namely CLIP4Clip [33], CLIP2Video [10] etc. These models are tested on three standard video retrieval benchmarks (§4.2) such as MSRVTT [52], MSVD [4] and DiDeMo [1].
@@
Our experiments (§5.1) reveal that objects & attributes are the most crucial to video retrieval followed by actions and syntax. Among video retrieval models, CLIP based models have a better compositional and syntactic understanding when compared with pretrained video models. We further perform detailed studies to fully judge how retrieval models perceive each of the components. We find that (§5.2) video retrieval models have a poor understanding of relationship between objects and its attributes. However, they are extremely sensitive to incorrect object references in the captions. Our studies on action understanding (§5.3) disclose that models have poor sense of action negation and replacing them with incorrect actions lead to slight decrease in video retrieval performance. Finally, we discover (§5.4) that models perform significantly better even without the right syntax. In summary, our contributions in this paper are as follows:
@@
  • Ours is the first work to comprehensively investigate the compositional and syntactic understanding of video retrieval models.
  • For this study, we propose a set of 10 tasks dealing with different aspects of compositionality and syntax.
@@ 2.1 Video retrieval
In a parallel direction, image features pre-trained on large amounts of image-text data have been adopted for the task of video retrieval. CLIP4Clip [33] is an end-to-end trainable video retrieval model based on CLIP [41] architecture in which frame features are extracted using clip image encoder and the temporal modelling is performed using a transformer encoder. A two-stage framework CLIP2Video [10] is proposed to enhance interaction among video features and video-text features for video-retrieval. Madasu et at. [36] used off-the-shelf multi-lingual data to enhance the performance of video-retrieval. All these video-retrieval models haven’t been tested for syntactic and compositional understanding. To the best of our knowledge, ours is the first work to comprehensively explore syntactic and compositional understanding of video retrieval models.
2.2 Syntactics
Transformer based language models [7, 20, 53] have achieved state-of-the-art results on most natural language understanding tasks [50, 51]. Hence, there has been a growing interest to explore the morphological capabilities of these models [55, 43, 49, 40, 14]. Since all the video retrieval models use pre-trained language models for encoding text captions, we build upon those works and investigate their syntactic understanding.
@@ 2.3 Compositionality
Although vision-language models pretrained on large amounts of data achieved state-of-the-results there has been a growing interest to understand the working of these models [27, 23, 6, 38, 23, 47]. These works mainly focus on compositional knowledge these models by proposing new benchmarks. Winoground [48] dataset was introduced in which a pair of text captions contain the same set of words but pertain to different images. The models are then tested for image and caption match. Another benchmark CREPE [13] was put forward to evaluate two aspects of compositionality: systematicity and productivity. This benchmark contains unseen compounds and atoms in the test split to evaluate the models’ generalization. Parcalabescu et al. [39] proposed VALSE dataset to measure visio-linguistic capabilities of pretrained vision and language models. AGQA-Decomp [13] is a new benchmark to measure compositional consistency for the task of Video Question Answering. All these works proposed new benchmarks for compositional reasoning in image-language models. Contrary to these, our work focuses on measuring compositionality of video retrieval models using the standard datasets and doesn’t require a new benchmark. Moreover, our experiments are evaluated on 12 models which are significantly higher than the frequency of models used in these works.
3 Compositional and Syntactic Understanding
@@
$Q_{synrem}$ Syntax removal guy wearing red shirt drives car talking
@@
In this section, we first define syntax and compositionality and subsequently establish the evaluation protocol for syntactic and compositional understanding in video retrieval models. For this evaluation, we augment the existing text captions and create new datasets that assess their syntactic and compositional understanding. We explain this protocol using an example test caption ($Q$) “a guy wearing a red shirt drives a car while talking” from the MSRVTT dataset. Table  1 summarizes different augmentation methods used for the proposed study.
@@ 3.1.1 Object & Attribute knowledge
Object & Attribute removal ($Q_{objattrrem}$): In this setup, we remove all the objects & attributes in the original caption $Q$ and the resulting caption is “wearing a drives a while talking”. Here guy, red shirt and car are the objects & attributes.
Object shift ($Q_{objshift}$): To test the VR models ability to relate objects with their attributes, we shift the places of objects in the captions. The modified caption is “a shirt wearing a red car drives a guy while talking”.
Object replacement ($Q_{objrep}$): We evaluate the VR models sensitivity to objects by randomly replacing the objects with an entirely different objects. The replaced caption is “a surf wearing a red mars drives a guy channel while talking”.
Object partial ($Q_{objpartial}$): In this setup, the VR models are given access to just 50% of the objects in the caption. This is to understand if the models perform any shortcuts while retrieving videos. Eg: “a wearing a red drives a car while talking”. Next, we introduce the tasks for evaluating action knowledge in VR models.
@@ 3.1.2 Action knowledge
Action removal ($Q_{actrrem}$): The actions present in the original captions are eliminated. The modified caption is “a red shirt a car while” as the actions wearing, drives are removed. This is to understand the influence of actions on the video retrieval performance.
Action negation ($Q_{actneg}$): A negation is added to all the actions in the captions resulting in the new caption “a guy not wearing a red shirt not drives a car while not talking” This tests the VR models ability to comprehend negation in the captions.
Action replacement ($Q_{actrep}$): In this setup, the actions are randomly replaced with a different set of actions. The replaced actions are neither antonyms nor synonyms. It checks if the models truly recognize the meaning of the action words. Next we present the evaluation protocol for syntactic understanding of VR models.
3.2 Syntactic understanding
In the previous section we elucidated the components for compositional reasoning in videos namely objects & attributes and actions. These components are bind together by syntax there by forming a meaningful caption. Let’s consider a part of the example described previously “a guy wearing a red shirt drives a car”, if the word “car” and “guy” are interchanged the resulting caption will be “a car wearing a red shirt drives a guy” which is not meaningful. Consequently, syntax also play a crucial role in video retrieval performance along with the compositonality. Subsequently, we put forward the evaluation protocol to measure syntactic understanding in video retrieval models.
Syntax removal ($Q_{synrem})$: Our first experiment focuses on the effect of syntax on VR models. We modify the caption by keeping just the objects & attributes, actions and eliminate any meaningful syntax among them. The resulting caption is “guy wearing red shirt drives car talking”.
Word order shuffle ($Q_{shuf})$: In this setup, all the words are shuffled in the caption. This destroys the order of compositionality and syntax. This tests the order sensitivity of VR models.
Word order reverse ($Q_{rev})$: In this setup, we preserve the word order except that in the reverse order. It evaluates the positional knowledge of video retrieval models. Next, we present the experiment set up for quantifying the compositional and syntactic understanding.
@@ 3.2 Syntactic understanding
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$
@@
Table 2: The table shows the results on MSRVTT [52] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{synrem}$ is the R@1 score on datasets that have excluded actions, attributes and syntax respectively.
@@
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$
@@
Table 3: The table shows the results on MSVD [4] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{synrem}$ is the R@1 score on datasets that have excluded actions, attributes and syntax respectively.
@@
Type Model Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$ Q $Q_{actrem}$ $Q_{objattrrem}$ $Q_{synrem}$
@@
Table 4: The table shows the results on DiDeMo [1] dataset in both text-to-video and video-to-text retrieval settings. $Q$ denotes the performance (R@1 score) on the original unchanged dataset. $Q_{actrem}$, $Q_{objattrrem}$ and $Q_{synrem}$ is the R@1 score on datasets that have excluded actions, attributes and syntax respectively.
@@ 4.3 Implementation
We use spacy [^1] https://spacy.io/usage/linguistic-features to identify parts-of-speech for all the words in the caption. We consider nouns, adverbs and adjectives as objects & attributes, verbs as actions and rest of the parts-of-speech as syntax. We use the exact set up used by the state-of-the-art video retrieval models and measure the performance on all the augmented datasets.
@@ 5 Results and Discussion
5.1 Objects & Attributes vs Actions vs Syntax: Do all of them matter?
@@
Our aim is to analyze the importance of three components: objects & attributes, actions and syntactics that make up a text query for retrieving videos. Hence, we test the video retrieval models with text captions that have missing objects & attributes ($Q_{objattrrem}$), actions ($Q_{actrem}$) and syntax ($Q_{synrem}$). Tables  2,  3 and  4 show the results on MSRVTT, MSVD and DiDeMo datasets respectively. It is evident from the table that there is a drop in video retrieval performance when tested with text captions that don’t have actions ($Q_{actrem}$). The drop is more pronounced among CLIP based models than pretrained video models. This shows that actions play a role for retrieving correct videos. However, we see that the performance drop is not as expected. Videos are time-series image frames which can have same attributes. In those scenarios, actions help in differentiating those videos. We see this effect when the R@1 is lower among pretrained video models and higher among CLIP based models. When the performance is lower, actions do not play a significant role in video retrieval and hence the videos can be retrieved without them in the text caption. On the contrary if R@1 score is higher, we see a notable decline. This is due to the robust video representations of CLIP based models as compared to pretrained video models. CLIP based models accurately encode video representations but, when the differentiating factor among videos i.e actions are missing in text captions leads to retrieval of incorrect videos. In short caption length datasets like MSRVTT and MSVD, we notice a significant drop in performance as compared with DiDeMo which is a paragraph ($>$ 1 sentences) dataset. This is because text captions in DiDeMo contains detailed description of the videos and hence, missing actions didn’t lead to drop in performance as compared to MSRVTT and MSVD. It demonstrates that actions are not essential in paragraph-video retrieval.
@@
Next, we analyze the performance of video retrieval models tested with text captions without syntactics ($Q_{synrem}$). From the table, it is clear that there is a reduction in R@1 without the syntax in the text captions. It validates that syntactics are necessary for retrieving correct ground truth videos. For MSRVTT, we observe that models tested without syntax under-perform compared to actions in the text captions and the average difference in performance is 2%. The reverse is true for MSVD and DiDeMo datasets where there is a huge difference of 9%. In addition, we also notice that CLIP based models are more sensitive to syntax than pretrained video models. Finally, we evaluate the video retrieval models on text captions in the absence of objects & attributes ($Q_{objattrrem}$). As seen from the results, these models perform poorly (a drop in 20%) which underscores the significance of objects & attributes. We also notice that $Q_{attrrem}$ trails $Q_{actrem}$ and $Q_{synrem}$ by a huge margin. This difference is more striking among CLIP based models as opposed to pretrained video models.